OpenStudy (anonymous):

WILL GIVE MEDAL! I NEED HELP WITH THESE QUICK): if f(x)=2x^2-4x and g(x)=5-2x, evaluate f(x)-g(x) for x=5 A. 15 B. 5 C. 35 D. 20 If f(x) = 5x – 3 and g(x) = 3x – 3, find f(x) – g(x). A. 2x B. 8x – 6 C. 2x – 6 D. 8x

OpenStudy (mathmale):

Line up the like terms of these two functions and then combine the like terms, remembering that you are subtracting g(x) from f(x). f(x) = 5x - 3 -g(x) = -3x -(-3) ----- = -------- = f(x) - g(x)

OpenStudy (anonymous):

you just plug in the 2x^2 -4x into f(x) and 5-2x into G(x) so it's (2x^2-4x)-(5-2x) or 2x^2-4x-5-(-2x) then you plug in 5 for all of the x's...

OpenStudy (anonymous):

so for the second question my answer would be C?

OpenStudy (mathmale):

Josh and I are proposing the same thing; I just find it helps some people to line up like terms in columns. Josh is right in that you plug in x=5 to obtain f(5)-g(5). Kristina: Please show your work...support your choice with data.

OpenStudy (anonymous):

all i did was subtracted 5-3 which equals 2 and 3-(-3) equals 6

OpenStudy (anonymous):

when f(x)=2x^2-4x, that's saying you can just insert 2x^2-4x when you see f(x) and the same with g(x) and 5-2x

OpenStudy (anonymous):

so f(x)-g(x) is really saying (2x^2-4x) - (5-2x)

OpenStudy (anonymous):

And when x=5 that means whenever you see an x you can just switch it with a 5. Does that help?

OpenStudy (anonymous):

so I would subtract 2x^2 from 5?

OpenStudy (mathmale):

K: f(x) = 5x - 3 -g(x) = -3x -(-3) ----- = -------- = f(x) - g(x) Finish this work, and then substitute 5 for x.

OpenStudy (anonymous):

when you put 5 in the place of x then 2x^2 will turn into 2*5^2 which is 2*25 or 50

OpenStudy (anonymous):

but for that problem I don't have to substitute anything @mathmale

OpenStudy (mathmale):

K: Have you read the problem statement carefully? if f(x)=2x^2-4x and g(x)=5-2x, evaluate f(x)-g(x) for x=5 Look at the last part of this sentence.

OpenStudy (anonymous):

you gave me different numbers above lol @mathmale

OpenStudy (anonymous):

okay so what would be the next step after we have 50 now? @Josh55

OpenStudy (mathmale):

Are you sure I did that? If so, please show me where I gave you "different numbers."

OpenStudy (anonymous):

can you see how you get 50?

OpenStudy (mathmale):

if f(x)=2x^2-4x and g(x)=5-2x, evaluate f(x)-g(x) for x=5 I suggested that you add up like terms as follows: f(x) = 5x - 3 -g(x) = -3x -(-3) ----- = -------- = f(x) - g(x) Finish this work, and then substitute 5 for x. Last time I checked, 5x-3x=2x, and so, if x=5, 2x=2(5) =10. and that's it. Where'd that 50 come from?

OpenStudy (mathmale):

I don't mean to abandon you, but unfortunately I have another commitment to attend to now. Good luck. See you both again on OpenStudy.

OpenStudy (anonymous):

the first problem is the one you substitute the second problem you do not. you keep giving me the second problem and saying i have to substitute @mathmale lol

OpenStudy (anonymous):

yes i do.. @Josh55

OpenStudy (anonymous):

ok :) now do the exact same thing in (2x^2-4x) - (5-2x)

OpenStudy (anonymous):

and that'll give you the answer

OpenStudy (anonymous):

2(5)^2-4(5)-(5-2(5) so just solve it like that?

OpenStudy (anonymous):

@Josh55

OpenStudy (anonymous):

and the answer would be 35 ?:)

OpenStudy (anonymous):

Yes, BUT be sure to ADD 2(5) at the very end because there's a double negative

OpenStudy (anonymous):

way to go

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

you're welcome, keep it up!